OpenStudy (anonymous):

Can anyone explain this? If you increase something by 40% and decrease the result by 40%, how much less %? The answer is 16% but I don't understand. Thanks!

Miracrown (miracrown):

Let something be x now increase that something by 40% This means x--> 1.4 x Now decrease that new something by 40% this means 1.4 x --> 0.6 (1.4x) = 0.84 x This means the final result: 0.84x is 16% (0.16) less than the original object "x". Make sense? :)

OpenStudy (anonymous):

Sorry, I'm still confused. Why is it 1.4 not 0.4? Isn't 40% = 4/10 = 0.4? Where did 0.6 come from?

Miracrown (miracrown):

Okay so when you increase an object by 40%. You take itself "1" and add "0.4" of itself to it. This means (1+0.4)*x =1.4x When you decrease by 40% you take away 0.4 of the object. (1-0.4)*y = 0.6*y

Miracrown (miracrown):

More concretely, say that you have 20 apples. I give you an additional 40% of apples. That means I give you 0.4*20 = 8 apples. So you now have 20 + 8 = 28 apples.

Miracrown (miracrown):

Now I take away 40% of your present amount of apples. 0.4 * 28 = 11.2 apples Bad example because it's not an even number

Miracrown (miracrown):

but you end up with 28-11.2 apples.

Miracrown (miracrown):

So the moral of the story is. Don't let anyone even give you a percent increase and then a percent decrease on your new amount! You lose

OpenStudy (anonymous):

Thank you so much, I understand now! :)